OpenStudy (anonymous):

IM VERY CONFUSED ON HOW TO DO THIS PLEASE HELP Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)=(x-7)/(x+3) and g(x)=(-3x-7)/(x-1)

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

to find f(g(x)) just replace x in f(x) with g(x) then do the algebra to see where you end up

OpenStudy (campbell_st):

so f(g(x)) is \[f(g(x)) = \frac{\frac{-3x -7}{x -1} - 7}{\frac{-3x -7}{x + 1} + 3}\]

OpenStudy (campbell_st):

oops I have thw wrong sign in the denominator it should be x - 1

OpenStudy (anonymous):

Refer to the attachment from Mathematica.

OpenStudy (anonymous):

haha yeah i just caught that. i understand that part i just dont know what to do after that

OpenStudy (campbell_st):

so I'd expect the next step is common denominators \[f(g(x)) = \frac{\frac{(-3x -7) - 7(x -1)}{x -1}}{\frac{(-3x -7) + 3(x -1)}{x -1}}\] then away you go

OpenStudy (anonymous):

oh ok thanks i think i got it from here

OpenStudy (campbell_st):

so just check.... my signs and things... but it is basically substitution, then getting common denominators and then simplifying and at least you are told the answer should be x. hope it helps